justinpm709:

After years of overhunting, environmental scientists have reintroduced wolverines to Yellowstone National Park. The initial number of wolverines reintroduced to the park was 88 and, after 5 years, the population is estimated to be around 578. Assuming an exponential growth pattern, what is the annual growth rate (rounded to the nearest tenth of a percent) of the new wolverine population in Yellowstone National Park? Hint: A(t) = A0(1 + r)t, where A(t) is the final amount, A0 is the initial amount, r is the growth rate expressed as a decimal amount, and t is time.

justjm:

Apply the general formula for exponential growth \( A(t) = A_{0}(1+r)^{t} \) You are given that \( A_0 \) is 88 and after \( t=5\) years, \(A(t)=578\) Plug that in \( 578 = 88(1+r)^5\) Solve for r \( \frac{578}{88} = (1+r)^5 \) r=0.4567 So the annual growth rate is 45.67%